OpenStudy (anonymous):

A painter charges $15.62 per hour, plus an additional amount for the supplies. If he made $173.87 on a job where he worked 6 hours, how much did the supplies cost? A:$93.72 B:$90.15 C:$80.15 D:$28.98

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (anonymous):

can someone help me

OpenStudy (anonymous):

15.62(t) +a = 173.87 t= no. of hours a= additional charges

OpenStudy (anonymous):

so a

OpenStudy (anonymous):

this is for study island

OpenStudy (anonymous):

@sangya21 are you thair

OpenStudy (anonymous):

15.62(6) + a = 173.87 a = 173.87 - 93.72

OpenStudy (anonymous):

its not a i got it worng

OpenStudy (anonymous):

Yeah. So what will be the correct answer?

OpenStudy (anonymous):

C I AM STUPID

OpenStudy (anonymous):

I AM SSSSOOO STUPIED

OpenStudy (anonymous):

You are not stupid. Just in hurry to solve prob

OpenStudy (anonymous):

thank you for the help is it ok if i can fan u

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

how old are you

OpenStudy (anonymous):

@sangya21 are you thair